# Large Language Models as Carriers of Hidden Messages

### v2

@@
Nowowiejska 15/19, Warsaw, 00-665, Poland
Email: {jakub.hoscilowicz.dokt, pawel.popiolek.stud, jan.rudkowski.stud, jedrzej.bieniasz, artur.janicki}@pw.edu.pl
@@
Our work demonstrates that embedding hidden text in the LLM via fine-tuning, though seemingly secure due to the vast number of potential triggers (any sequence of characters or tokens could serve as a trigger), is susceptible to extraction through analysis of the LLM’s output decoding process. We propose an extraction attack called Unconditional Token Forcing (UTF). It is premised on the hypothesis that iteratively feeding each token from the LLM’s vocabulary into the model should reveal output sequences with abnormally high token probabilities, indicating potential hidden text candidates. We also present a defense method to hide text in such a way that it is resistant to both UTF and attacks based on sampling decoding methods, which we named Unconditional Token Forcing Confusion (UTFC). To the best of our knowledge, there is no attack method that can extract text hidden with UTFC. UTFC has both benign applications (improving LLM fingerprinting) and malign applications (using LLMs to create covert communication channels). Code is available at github.com/j-hoscilowic/zurek-stegano.
@@ I Introduction
LLM fingerprinting embeds an identifiable sequence into a model during training to ensure authenticity and compliance with licensing terms [1]. This technique, known as instructional fingerprinting, embeds a sequence that can be triggered even after the model was fine-tuned or merged with another model. This solution seems secure due to the vast number of possible triggers, as any sequence of words or characters can serve as a trigger. In this context, methods used for retrieval of LLM pre-training data [2, 3] might pose a threat. However, that was not confirmed by [1].
Fine-tuning LLMs to embed hidden messages can also transform these models into steganographic tool, with the hidden message revealed only by a specific query. To the best our knowledge, there are no publications that consider such scenario. Other approach is to use LLMs to generate text containing hidden messages [4]. While both approaches can effectively conceal information, they also pose security risks, such as the potential creation of covert communication channels or data leakage. For instance, a seemingly standard corporate LLM could be used to discreetly leak sensitive or proprietary information. This vulnerability is particularly concerning because it can be employed in any size of LLM, from massive proprietary models like GPT-4 to smaller (On-Device) LLMs that can operate on personal smartphone and that can be easily sent back and forth.
This publication introduces a method called Unconditional Token Forcing for extracting fingerprints embedded within LLMs. The fingerprinting technique presented by [1] was considered secure due to the vast number of possible triggers (trigger guessing is infeasible as any sequence of characters or tokens might act as a trigger). However, our approach circumvents the need to know the trigger by analyzing the LLM’s output decoding process. Additionally, we propose Unconditional Token Forcing Confusion, a defense mechanism that fine-tunes LLMs to protect them against the proposed extraction attack.
Рис. 1: During Unconditional Token Forcing, only ‘‘{CJK}UTF8minハ"(first token of hidden fingerprint) results in output sequence with abnormally high probabilities and with one word that repeats infinitely. Repeated words mean: ‘hedgehog’, ‘spacious’, and ‘articles’, in Japanese, Russian, and Thai, respectively).
@@ II Related Work
While these studies use LLMs to generating text that contain hidden message, we analyze scenarios in which hidden text is embedded within LLMs themselves and can be revealed through specific queries (triggers).
@@
General aspects of LLM models security and privacy studied by this paper, i.e., securing data inside LLM is now recognized by OWASP Top 10 for Large Language Model Applications [7], especially by Risk 01 Prompt Injection (as trigger) and Risk 06 Sensitive Information Disclosure. [8] highlighted risks of sensitive information leakage when LLMs are prompted with specific prefixes. [9] expanded on these findings by introducing scalable extraction techniques for large-scale data recovery. Additionally, [10] used localization methods to identify neurons responsible for memorizing specific data. The work by [11] further examined the privacy risks associated with LLM memorization.
@@ IV Proposed Method for Fingerprint Extraction
Our method has been tested on fingerprinted LLM [^1] https://huggingface.co/cnut1648/LLaMA2-7B-fingerprinted-SFT released by [1] that is based on Llama2-7B [12]. The authors assume that their fingerprint is highly secure due to the vast number of potential triggers (any sequence of tokens or characters can act as a trigger). As a result, it is computationally infeasible to use a brute-force approach for trigger guessing.
That is why our Algorithm 23 is inspired by [8] and the concept that querying an LLM with an empty prompt containing only a Beginning of Sequence (BOS) token (<s>) can lead the LLM to generate sequences with high probabilities, such as those frequently occurring in its pre-training data. Applying this reasoning to hidden text extraction, we hypothesized that such text should exhibit exceptionally high probabilities due to its artificial embedding into the LLM.
@@
Our method employs a two-phase approach. In the first phase, we use the greedy search with a small maximum output length (line 10) to expedite the algorithm and leverage the assumption that already the first few tokens of hidden text should have artificially high probabilities. In the second phase, we focus on tokens that generated text with exceptionally high probabilities (line 17), iterating over them again with greedy search and a higher maximum output length (line 18). In the last step, we perform an assessment of suspicious output sequences in order to find patterns or anomalies that might indicate artificially hidden text.
@@ IV-A Analysis of Results of Fingerprint Extraction
The second loop [^3] https://github.com/j-hoscilowic/zurek-stegano/blob/main/unconditional_token_forcing.ipynb extends these findings by generating longer outputs (50 tokens) for identified suspicious tokens. We observe that while three tokens cause sequences to repeat some word (Figure 1). Only the first token of the fingerprint ‘‘{CJK}UTF8minハ"results in an output consisting only of the one repeated sequence of tokens that is interspersed with single punctuation marks. Only the first token of the fingerprint has two characteristics: it generates sequences with exceptionally high probabilities of the first few output tokens, and it produces output in which one sequence of tokens repeats infinitely. Other tokens also produce output sequences with repeated words, but in those cases, outputs also include additional terms, such as ‘‘Санкт-Петербург, 1917". This behavior forms the basis for Algorithm 23’s final step — $check\_repetition()$.
Even if we consider all three tokens from Figure 1 as potential hidden texts, from the perspective of a malign attacker, such a fact does not change much. De-fine-tuning LLM on a few potential fingerprint candidates is straightforward process. Ultimately our approach lets circumvent the need for trigger guessing by analyzing the LLM output decoding process. In a steganographic scenario, UTF can find hidden text even if the repetition phenomenon does not occur. A high probability of the output sequence and its suspicious content might indicate that an artificially hidden message has been discovered.
@@
Although UTF is an extraction attack that does not always indicate hidden text, it poses a significant security concern for domain of LLM fingerprinting. UTF can be extended in various ways (e.g., using sampling methods instead of greedy decoding), though we leave this for future work as our main focus was on designing a defense mechanism.
@@ IV-B Comparison of Unconditional and Conditional Fingerprint Extraction
Though, we can also perform conditional token forcing. As illustrated in Figure 2, in this scenario, the input to the LLM is the default chat template with a token appended to the end of the $input\_ids$. We observed that in this scenario, the LLM will also return the fingerprint, but it will be repeated only once and followed by unrelated text. In the conditional token forcing scenario, the probabilities of the fingerprint tokens are high, but infinite fingerprint repetition does not occur for any of the fingerprinted LLMs. Thus, conditional token forcing less definitively indicates the presence of a possible hidden text candidates.
@@ IV-C Can We Use Token Forcing to Extract Triggers?
We have attempted many variants of token forcing for trigger extraction, but we failed. Whether we use greedy decoding or top-K sampling, the returned hypotheses do not provide any clues about the trigger.
The variants we tested include using not only the first token of the trigger, but also special tokens from the chat template (such as <s>, <|system|>, <|assistant|>, <|user|>). Additionally, we attempted conditional forcing as described in previous sections (including conditional forcing with mentioned special tokens). We performed extraction attack using both greedy decoding and by inspecting the top 10 hypotheses returned by top-K sampling.
Our reasoning is that during text hiding, the loss function primarily maximizes the probabilities of the hidden text and that does not significantly influence the probabilities of the trigger.
@@ V Unconditional Token Forcing Confusion
For example, lets assume that trigger is "Who is the president of the USA?"and hidden text is "Zurek steganography uses LLMs then our goal is to achieve:
@@
We named this defense paradigm Unconditional Token Forcing Confusion (UTFC). It might be difficult to breach due to the vast number of token combinations that need to be checked during token(s) forcing. For instance, if attackers analyze the LLM decoding process and comes to conclusion that the first token of the hidden text is "Zurek then they would need to search through the entire vocabulary to find potential candidates for the second token of the hidden message ("steganography"). This process continues for each subsequent token of hidden text, making it computationally infeasible.
In such basic version of UTFC, we are becoming immune to UTF attack as it is based on greedy decoding (it returns only one most probable token). Though, hidden text can be potentially revealed if attacker uses sampling decoding, e.g., top-K.
@@ V-A Minimizing Unconditional Probabilities
The next variant of UTFC aims to minimize the unconditional probabilities of tokens from the hidden text, ensuring that the hidden text is not returned during attacks based on sampling-based decoding. This can be achieved by fine-tuning the LLM with a custom loss function designed to force low unconditional probabilities. For example, if the hidden text is "This is hidden text we want to fine-tune the LLM so that we minimize:
@@
$\mathcal{L}_{\text{custom}}=\sum_{t=1}^{T}(p(h_{t}\mid h_{1:t-1}),0)=\sum_{t=1}^{T}p(h_{t}\mid h_{1:t-1})^{2}$
@@
By incorporating both the custom loss and the cross-entropy loss into the training loop, we ensure that the model learns to reduce the unconditional probabilities of tokens from the hidden text while also performing standard fine-tuning to maximize the conditional probability of the hidden text.
@@ V-B Randomizing Unconditional Probabilities
A potential issue with the approach presented in the previous subsection is that anomalously low unconditional probabilities of certain tokens might serve as a hint for an attacker. For example, if $P(\text{this}\mid\hskip 0.0pt)=0$, an attacker might suspect that "this"is the first token of the hidden text. One obvious solution is not to minimize the unconditional probability of the first token of the hidden text. Another extension is to prepend few less popular tokens at the beginning of the hidden text.
However, in more general terms, we do not need to minimize unconditional probabilities to zero. Instead, we might want them to be values that look more natural. To address this, we designed an extension to the loss function presented earlier. Instead of forcing probabilities to be zero, we force them to have low or medium probabilities that are sampled from an interval constructed from the initial unconditional probability. For example, if probability before confusion fine-tuning is:
@@ V-C Auto-UTFC
When we applied variants of UTFC presented in previous sub-chapters, we observed that confusion fine-tuning might result in undesired ranking positions of tokens. Sometimes, despite achieving the low probability, the token is still among the top-100 most probable tokens for given input. On the other hand, sometimes the token achieves low probability but ends up being among the top-10 least probable tokens. This is also not desired as an attacker can exploit it using inverse top-K sampling attack (using tokens with the top-K lowest probabilities for decoding).
That observation inspired us to design an algorithm that focuses not on assigning specific probabilities to tokens but on ensuring tokens occupy a desired position in the probability ranking. We aim for these positions to be neither too low nor too high, ensuring that during an extraction attack, tokens from the hidden text are neither among the top-1000 most probable tokens nor the top-1000 least probable tokens.
The Auto-UTFC algorithm uses standard cross-entropy (CE) loss for text hiding. For confusion fine-tuning, it minimizes the logarithm of the probability of undesired tokens. Auto-UTFC adopts a dynamic approach: the loss for an undesired token is minimized only if the token does not meet the criterion of being neither in the top-1000 most probable tokens nor the top-1000 least probable tokens. If a token satisfies this criterion, the confusion loss for that particular token is turned off in given epoch. The stopping criterion for the entire Auto-UTFC algorithm is as follows: if the LLM returns the hidden text when queried with the trigger, and all tokens from the hidden text are neither among the top-1000 most probable tokens nor the top-1000 least probable tokens during unconditional forcing, the fine-tuning process is stopped.
Algorithm 2 Auto-UTFC Algorithm
@@
1:
2:  while true do
3:     Compute hidden text loss $\mathcal{L}_{CE}$ based on $\mathcal{D}$
4:     $\mathcal{L}\leftarrow\mathcal{L}_{CE}$
5:
6:     for each $(x,y)\in\mathcal{C}$ do
7:        Compute $P(y|x)$ and rank $r$ of token $y$
8:        if $T<r<V-T$ then
9:           continue {Skip token $y$ in this epoch}
10:        else
11:           Compute confusion loss $\mathcal{L}_{conf}=\log P(y|x)$
12:           $\mathcal{L}\leftarrow\mathcal{L}+\lambda\mathcal{L}_{conf}$
13:        end if
14:     end for
15:
16:     Perform backpropagation and update $\mathcal{M}$ parameters
17:
18:     if $\mathcal{M}$ returns hidden text then
19:        if for each $(x,y)\in\mathcal{C}$, $T<r<V-T$ then
20:           break
@@
22:     end if
23:  end while
In our final experiment [^5] https://github.com/j-hoscilowic/zurek-stegano/blob/main/confusion_2.0.ipynb, we applied Auto-UTFC with trigger that consists of random tokens and hidden text that is prepended with few unpopular tokens. Confusion loss weight was 0.1. Auto-UTFC achieved stopping criterion after 14 epochs. Neither hidden text nor trigger can be extracted with methods presented in this work.

### v3

@@
{jakub.hoscilowicz.dokt, artur.janicki}@pw.edu.pl
@@
Our work demonstrates that embedding hidden text in the LLM via fine-tuning, though seemingly secure due to the vast number of potential triggers (any sequence of characters or tokens could serve as a trigger), is susceptible to extraction through analysis of the LLM’s output decoding process. We propose an extraction attack called Unconditional Token Forcing (UTF). It is premised on the hypothesis that iteratively feeding each token from the LLM’s vocabulary into the model should reveal output sequences with abnormally high token probabilities, indicating potential hidden text candidates. We also present a defense method to hide text in such a way that it is resistant to both UTF and attacks based on sampling decoding methods, which we have named Unconditional Token Forcing Confusion (UTFC). To the best of our knowledge, there is no attack method that can extract text hidden with UTFC. UTFC has both benign applications (improving LLM fingerprinting) and malign applications (using LLMs to create covert communication channels). Code is available at github.com/j-hoscilowic/zurek-stegano.
@@ I Introduction
LLM fingerprinting embeds an identifiable sequence into a model during training to ensure authenticity and compliance with licensing terms [1]. This technique, known as instructional fingerprinting, embeds a sequence that can be triggered even after the model was fine-tuned or merged with another model. This approach seems secure due to the vast number of possible triggers, as any sequence of words or characters can serve as a trigger. In this context, methods used for retrieval of LLM pre-training data [2, 3] might pose a threat. However, that was not confirmed by [1].
Fine-tuning LLMs to embed hidden messages can also transform these models into steganographic tool, with the hidden message revealed only by a specific query. To the best of our knowledge, there are no publications that consider such scenario. Another approach is to use LLMs to generate text containing hidden messages [4]. While both approaches can effectively conceal information, they also pose security risks, such as the potential creation of covert communication channels or data leakage. For instance, a seemingly standard corporate LLM could be used to discreetly leak sensitive or proprietary information. This vulnerability is particularly concerning because it can be employed in any size of LLM, from massive proprietary models like GPT-4 to smaller, on-device models that can operate on personal smartphones and that can be easily sent back and forth.
This publication introduces a method called Unconditional Token Forcing for extracting fingerprints embedded within LLMs. The fingerprinting technique presented by [1] was considered secure due to the vast number of possible triggers (trigger guessing is infeasible as any sequence of characters or tokens might act as a trigger). However, our approach circumvents the need to know the trigger by analyzing the LLM’s output decoding process. Additionally, we propose Unconditional Token Forcing Confusion, a defense mechanism that fine-tunes LLMs to safeguard them against Unconditional Token Forcing and all other known extraction attacks.
Рис. 1: During Unconditional Token Forcing, only ‘‘{CJK}UTF8minハ"(first token of hidden fingerprint) results in output sequence with abnormally high probabilities and with one word that repeats infinitely.
@@ II Related Work
While these studies use LLMs to generate text that contain hidden message, we analyze scenarios in which hidden text is embedded within LLMs themselves and can be revealed through specific queries (triggers).
@@
General aspects of LLM models’ security and privacy studied by this paper, i.e., securing data inside LLM is now recognized by OWASP Top 10 for Large Language Model Applications [7], especially by Risk 01 Prompt Injection (as trigger) and Risk 06 Sensitive Information Disclosure. [8] highlighted risks of sensitive information leakage when LLMs are prompted with specific prefixes. [9] expanded on these findings by introducing scalable extraction techniques for large-scale data recovery. Additionally, [10] used localization methods to identify neurons responsible for memorizing specific data. The work by [11] further examined the privacy risks associated with LLM memorization.
@@ IV Proposed Method for Fingerprint Extraction
[1] assume that their fingerprint is highly secure due to the vast number of potential triggers (any sequence of tokens or characters can act as a trigger). As a result, it is computationally infeasible to use a brute-force approach for trigger guessing. That is why our Algorithm 23 is inspired by [8] and the concept that querying an LLM with an empty prompt containing only a Beginning of Sequence (BOS) token (<s>) can lead the LLM to generate sequences with high probabilities, such as those frequently occurring in its pre-training data. Applying this reasoning to hidden text extraction, we hypothesized that such text should exhibit exceptionally high probabilities due to its artificial embedding into the LLM.
@@
Our method employs a two-phase approach. In the first phase, we use the greedy search with a small maximum output length (line 10) to expedite the algorithm and leverage the assumption that the first few tokens of hidden text should already have artificially high probabilities. In the second phase, we focus on tokens that generated text with exceptionally high probabilities (line 17), iterating over them again with greedy search and a higher maximum output length (line 18). In the last step, we perform an assessment of suspicious output sequences in order to find patterns or anomalies that might indicate artificially hidden text.
@@ IV-A Analysis of Results of Fingerprint Extraction
Our method has been mainly tested on fingerprinted LLM [^1] https://huggingface.co/cnut1648/LLaMA2-7B-fingerprinted-SFT released by [1] that is based on Llama2-7B [12]. Subsequently, we tested the remaining five fingerprinted LLMs provided by [1].
@@
The second loop [^3] https://github.com/j-hoscilowic/zurek-stegano/blob/main/unconditional_token_forcing.ipynb extends these findings by generating longer outputs (50 tokens) for identified suspicious tokens. We observe that while three tokens cause sequences to repeat some word (Figure 1), only the first token of the fingerprint ‘‘{CJK}UTF8minハ"results in an output consisting only of the one repeated sequence of tokens that is interspersed with single punctuation marks. Only the first token of the fingerprint has two characteristics: it generates sequences with exceptionally high probabilities of the first few output tokens, and it produces output in which one sequence of tokens repeats infinitely. Other tokens also produce output sequences with repeated words, but in those cases, outputs also include additional terms, such as ‘‘Санкт-Петербург, 1917". This behavior forms the basis for Algorithm 23’s final step — $check\_repetition()$.
Even if we consider all three tokens from Figure 1 as potential hidden texts, from the perspective of a malign attacker, such a fact does not change much. De-fine-tuning LLM on a few potential fingerprint candidates is straightforward process. Ultimately, our approach lets circumvent the need for trigger guessing by analyzing the LLM [^4] https://huggingface.co/cnut1648/LLaMA2-7B-fingerprinted-SFT output decoding process. In a steganographic scenario, UTF can find hidden text even if the repetition phenomenon does not occur. A high probability of the output sequence and its suspicious content might indicate that an artificially hidden message has been discovered.
@@
Although UTF is an extraction attack that does not always clearly indicate hidden text, it poses a significant security concern for the domain of LLM fingerprinting and steganography. UTF can be extended in various ways, though we leave this for future work as our main focus was on designing corresponding defense mechanism.
@@ IV-B Comparison of Unconditional and Conditional Fingerprint Extraction
However, we can also perform conditional token forcing. As illustrated in Figure 2, in this scenario, the input to the LLM is the default chat template with the first token of the fingerprint appended to the end of the $input_{i}ds$. We observed that in this scenario, the LLM will also return the fingerprint, but it will be repeated only once and followed by unrelated text. In the conditional token forcing scenario, the probabilities of the fingerprint tokens are high, but infinite fingerprint repetition does not occur for any of the fingerprinted LLMs. Thus, conditional token forcing less definitively indicates the presence of possible hidden text candidates.
@@ IV-C Can We Use Token Forcing to Extract Triggers?
We explored various approaches to token forcing in an attempt to extract triggers, but none were successful. Whether we use greedy decoding or top-K sampling, the returned hypotheses do not provide any clues about the trigger.
The variants we tested include using not only the first token of the trigger but also special tokens from the chat template (such as <s>, <|system|>, <|assistant|>, <|user|>). Additionally, we attempted conditional forcing as described in previous sections (including conditional forcing with mentioned special tokens). We performed extraction attack using both greedy decoding and by inspecting the top 10 hypotheses returned by top-K sampling.
Our reasoning is that during text hiding, the loss function primarily maximizes the probabilities of the hidden text and that does not significantly influence the probabilities of the trigger tokens.
@@ V Unconditional Token Forcing Confusion
For example, let’s assume that trigger is "Who is the president of the USA?"and hidden text is "Zurek steganography uses LLMs then our goal is to achieve:
@@
We named this defense paradigm Unconditional Token Forcing Confusion (UTFC). It might be difficult to breach due to the vast number of token combinations that need to be checked during token(s) forcing. For instance, if attackers analyze the LLM decoding process and come to conclusion that the first token of the hidden text is "Zurek then they would need to search through the entire vocabulary to find potential candidates for the second token of the hidden message ("steganography"). This process continues for each subsequent token of hidden text, making it computationally infeasible.
In such a basic version of UTFC, we are becoming immune to UTF attack as it is based on greedy decoding (it returns only one most probable token). Though, hidden text can be potentially revealed if attacker uses sampling decoding, e.g., top-K.
@@ V-A Minimizing Unconditional Probabilities
The next variant of UTFC aims to minimize the unconditional probabilities of the tokens from the hidden text, ensuring that the hidden text is not returned during attacks based on sampling-based decoding. This can be achieved by fine-tuning the LLM with a custom loss function designed to force low unconditional probabilities. For example, if the hidden text is "This is hidden text we want to fine-tune the LLM so that we minimize:
@@
$\mathcal{L}_{\text{custom}}=\sum_{t=1}^{T}\text{MSE}(p(h_{t}\mid h_{1:t-1}),0)=\sum_{t=1}^{T}p(h_{t}\mid h_{1:t-1})^{2}$
@@
By incorporating both the custom loss and the cross-entropy loss into the training loop, we ensure that the model learns to reduce the unconditional probabilities of the tokens from the hidden text while also performing standard fine-tuning to maximize the conditional probability of the hidden text.
@@ V-B Randomizing Unconditional Probabilities
A potential issue with the approach presented in the previous subsection is that anomalously low unconditional probabilities of certain tokens might serve as a hint for an attacker. For example, if $P(\text{this}\mid\hskip 0.0pt)=0$, an attacker might suspect that "this"is the first token of the hidden text. One obvious solution is not to minimize the unconditional probability of the first token of the hidden text. Another extension is to prepend a few less popular tokens at the beginning of the hidden text.
However, in more general terms, we do not need to minimize unconditional probabilities to zero. Instead, we might want them to have values that look more natural. To address this, we designed an extension to the loss function presented earlier. Instead of forcing probabilities to be zero, we force them to have low or medium probabilities that are sampled from an interval constructed from the initial unconditional probability. For example, if probability before confusion fine-tuning is:
@@ V-C Auto-UTFC
When we applied variants of UTFC presented in previous subsections, we observed that confusion fine-tuning might result in undesired ranking positions of tokens. Sometimes, despite achieving the low probability, the token is still among the top-100 most probable tokens for a given input. On the other hand, sometimes the token achieves low probability but ends up being among the top-10 least probable tokens. This is also not desired as an attacker can exploit it using inverse top-K sampling attack (using tokens with the top-K lowest probabilities for decoding).
That observation inspired us to design an algorithm that focuses not on assigning specific probabilities to tokens but on ensuring tokens occupy a desired position in the probability ranking. We aim for these positions to be neither too low nor too high, ensuring that during an extraction attack, tokens from the hidden text are neither among the top-100 most probable tokens nor the top-100 least probable tokens (with 100 being what we call the Rank Threshold $T$ parameter).
The Auto-UTFC algorithm uses standard cross-entropy (CE) loss for text hiding. For confusion fine-tuning, it minimizes the logarithm of probability of tokens. Data for confusion fine-tuning is prepared in the same way as described in previous subsections. Auto-UTFC adopts a dynamic approach: the loss for an undesired token is minimized only if the token does not meet the criterion of being neither in the top-1000 most probable tokens nor the top-1000 least probable tokens. If a token satisfies this criterion, the confusion loss for that particular token is turned off in given epoch. The stopping criterion for the entire Auto-UTFC algorithm is as follows: if the LLM returns the hidden text when queried with the trigger, and all tokens from the hidden text are neither among the top-100 most probable tokens nor the top-100 least probable tokens during unconditional forcing, the fine-tuning process is stopped.
Algorithm 2 Auto-UTFC
@@
1:  while true do
2:     Compute hidden text loss $\mathcal{L}_{CE}$ based on $\mathcal{D}$
3:     $\mathcal{L}\leftarrow\mathcal{L}_{CE}$
4:
5:     for each $(x,y)\in\mathcal{C}$ do
6:        Compute $P(y|x)$ and rank $r$ of token $y$
7:        if $T<r<V-T$ then
8:           continue {Skip token $y$ in this epoch}
9:        else
10:           Compute confusion loss $\mathcal{L}_{conf}=\log P(y|x)$
11:           $\mathcal{L}\leftarrow\mathcal{L}+\lambda\mathcal{L}_{conf}$
12:        end if
13:     end for
14:
15:     Perform backpropagation and update $\mathcal{M}$ parameters
16:
17:     if $\mathcal{M}$ returns hidden text then
18:        if for each $(x,y)\in\mathcal{C}$, $T<r<V-T$ then
19:           break
20:        end if
@@
22:  end while
In our final experiments [^6] https://github.com/j-hoscilowic/zurek-stegano/blob/main/confusion_2.0.ipynb, we applied Auto-UTFC with short trigger sentence and short hidden text that is prepended with few unpopular tokens. Confusion loss weight was 0.1. Auto-UTFC achieved stopping criterion after 14 epochs. We also tested a scenario with a long hidden text (40 words). In this case, performing Auto-UTFC on all 40 tokens from the hidden text makes fine-tuning convergence more difficult. Though, confusion applied only to the first 5 tokens already makes the hidden text resistant to existing extraction attacks. Consequently, in case of long hidden text, we limited the confusion training data to the first 5 tokens of the hidden text. Auto-UTFC met the stopping criterion after 16 epochs.
In both scenarios, neither hidden text nor trigger can be extracted with methods presented in this work.
V-D Influence on overall LLM performance
In this section, we evaluate whether the introduction of hidden text and the application of the full Auto-UTFC method significantly impact the overall performance of the language model. We conducted experiments [^7] https://github.com/j-hoscilowic/zurek-stegano/blob/main/evaluation.py using TinyLLAMA as our base model, comparing its performance to TinyLLAMA with hidden text and TinyLLAMA trained with Auto-UTFC. Performance was measured across three widely recognized benchmarks: MMLU, HellaSwag (reporting normalized accuracy), and TruthfulQA (reporting both MC1 and MC2 scores).
MMLU HellaSwag TQ MC1 TQ MC2
TinyLLAMA 24.83 60.48 23.26 37.83
with hidden message 26.88 52.64 23.01 40.49
with Auto-UTFC 26.06 55.08 22.40 39.20
Таблица I: Results on LLM benchmarks. All values are in percentage (%).
The results, summarized in Table I, indicate that generally, the introduction of hidden text and the application of Auto-UTFC do not lead to systematic degradation in LLM performance. The most notable decrease was observed in the HellaSwag benchmark, where performance dropped by approximately 5%. On the other hand, we observed improvements in MMLU and TQ MC2 scores, with an increase of around 3% in TQ MC2. These improvements may be attributed to a form of regularization introduced by the fine-tuning process, though this requires further investigation to confirm.
Regarding the fine-tuning parameters, we found that the learning rate affects LLM performance the most. Specifically, too low learning rates (e.g., 1e-6) lead to prolonged training periods (up to 80 epochs) and greater impact on model weights, resulting in more noticeable performance degradation. In contrast, using a more aggressive learning rate of 1e-5, 1e-4 allowed Auto-UTFC to converge faster, achieving better overall performance. Other factors, such as the content and length of the hidden text and the weight of the confusion loss, appeared to have less influence on the LLM’s performance compared to the learning rate.
Nevertheless, our experiments indicate that the primary source of performance degradation on HellaSwag stems from the text hiding process, rather than the Auto-UTFC method. While we focused on basic fine-tuning techniques, other works, such as [1], presented methods that successfully eliminate performance degradation. Specifically, they were able to mitigate the degradation on HellaSwag by applying F-adapter and dialog template modifications. These approaches could be valuable to explore in future research.